OpenStudy (anonymous):

Calib received $50 for his birthday that he put in his piggy bank. Each week, he puts 2 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 2w + 50 where w is the number of weeks. His sister Lara makes a table of reasonable values for w to determine how much money Calib will eventually have. If one of the values for w Lara selects is 44, what would be her corresponding entry for D? A: 138 B: 96 C:77 D: 294

OpenStudy (wolf1728):

Dollars = 2 * weeks + 50 For w = 44 Dollars = 2 * 44 + 50 Dollars = 88 + 50 Dollars = $138 Answer is A

OpenStudy (anonymous):

Thank you so much! this helped big time! do you think you can help me with 1 more?